iler Manufacturing has 4 million shares of common stock outstanding. The current share price is $76, and the book value per share is $5. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $90 million, has a 5 percent coupon, and sells for 94 percent of par. The second issue has a face value of $70 million, has a 6 percent coupon, and sells for 104 percent of par. The first issue matures in 20 years, the second in 3 years. The most recent dividend was $4.8 and the dividend growth rate is 8 percent. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 35 percent. Required: What is the company's WACC? (Do not round your intermediate calculations.)
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=76*4000000
=304000000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*90000*0.94
=84600000
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=1000*70000*1.04
=72800000
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=304000000+84600000+72800000
=461400000
Weight of equity = MV of Equity/MV of firm
Weight of equity = 304000000/461400000
W(E)=0.6589
Weight of debt = MV of Bond/MV of firm
Weight of debt = 157400000/461400000
W(D)=0.3411
Cost of equity
As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
76 = 4.8 * (1+0.08) / (Cost of equity - 0.08)
Cost of equity% = 14.82
Cost of debt
Bond1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =20x2
940 =∑ [(5*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^20x2
                   k=1
YTM1 = 5.4983086304
Bond2
                  K = Nx2
Bond Price   =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                   K =3x2
1040 =∑ [(6*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^3x2
                    k=1
YTM2 = 4.56
Firm cost of debt=YTM1*(MV bond1)/(MV bond1+MV bond2)+YTM2*(MV bond2)/(MV bond1+MV bond2)
Firm cost of debt=5.4983086304*(84600000)/(84600000+72800000)+4.56*(84600000)/(84600000+72800000)
Firm cost of debt=5.06%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 5.06*(1-0.35)
= 3.289
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=3.29*0.3411+14.82*0.6589
WACC =10.89%